A student randomly selects 16 DVD’s at a store. The prices of the DVD’s have a standard deviation of $2.25.

Construct a 95% confidence interval for the population standard deviation. Assume the data are normally distributed.
Answer #1

Solution :

Given that,

s = 2.25

s2 = 5.0625

\chi2R = \chi 2\alpha/2,df = 27.488

\chi2L = \chi 21 - \alpha /2,df = 6.262

The 95% confidence interval for \sigma is,

\sqrt(n - 1)s2 / \chi 2\alpha/2 < \sigma < \sqrt (n - 1)s2 / \chi 21 - \alpha /2

\sqrt(15)(5.0625) / 27.488 < \sigma < \sqrt (15)(5.0625) / 6.262

1.66 < \sigma < 3.48

(1.66 , 3.48)